OpenStudy (anonymous):

Find the integral of cos^2x dx.please help

OpenStudy (lgbasallote):

[-sin^2(x)]/2

OpenStudy (lgbasallote):

it's a combination of the power rule thingy and the sinx = cosxdx thingy

OpenStudy (perl):

and the other thingy

OpenStudy (mimi_x3):

\[\int\limits \cos^{2}x = \frac{1}{2}(1+\cos2x)dx = \frac{1}{2} (x+\frac{1}{2}\sin2x) +c\]

OpenStudy (lgbasallote):

there are a lot of thingies here, aren't there? :))

OpenStudy (anonymous):

Mimi,i understand your steps but not the next one,my final answer is 1\4sin(2x)+1/2x+c which is wrong,it si supposed to be 1\4sin(2x)+2x+c.how do they get 2x?

OpenStudy (lgbasallote):

wolframalpha said something like 1.2(x +sinxcosx) + C haha so embarassing to admit i'm wrong..but i am :))

OpenStudy (lgbasallote):

that's 1/2

OpenStudy (turingtest):

mimi is right, so it seems your book is wrong

OpenStudy (anonymous):

you can say that again T,@ all-thax for ur help.Igbasallote dont be embarrased one who makes no mistakes never do anything